OpenStudy (anonymous):

Given the function g(x) = -4x + 5, compare and contrast g(2) and g(-4). Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(-4). The value of g(2) is smaller than the value of g(-4). The value of g(2) is the same as the value of g(-4). The values of g(2) and g(-4) cannot be compared.

OpenStudy (anonymous):

@texaschic101 @hartnn @ganeshie8

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Well all you would do is plug in 2 for 'x' and then do the same thing by plugging in -4 for 'x' and see which comes out bigger compare to the given choices...

OpenStudy (anonymous):

A

OpenStudy (johnweldon1993):

Try again, remember to watch the negative values when doing the -4

OpenStudy (anonymous):

C :)

OpenStudy (mathmale):

Jadee: Consider explaining your choice(s) of A and/or C. I'd like to know your reasoning, so that I can give you appropriate feedback.

OpenStudy (johnweldon1993):

Here we'll do it out \[\large g(x) = -4x + 5\] \(\large g(2)\) means replace every 'x' value we have with 2 and \(\large g(-4)\) means replace every 'x' value with -4 so \[\large g(2) = -4(2) + 5\] what does that come out to?

OpenStudy (anonymous):

-3

OpenStudy (johnweldon1993):

Correct, now for the next one \[\large g(-4) = -4(-4) + 5\] what does that come out to?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

so they cant be compared

OpenStudy (johnweldon1993):

Not quite \[\large g(-4) = -4(-4) + 5\] \[\large g(-4) = 16 + 5\] \[\large g(-4) = 21\] does that make sense? since -4 times -4 is 16 *negative times negative = positive*

OpenStudy (anonymous):

yes it makes sense

OpenStudy (johnweldon1993):

Alright so we have \[\large g(2) = -3\] \[\large g(-4) = 21\] What conclusion can you make with that information?

OpenStudy (anonymous):

the value of 2 is smaller

OpenStudy (johnweldon1993):

Correct, which equates to answer choice...?

OpenStudy (mathmale):

Choose the statement that is true concerning these two values. The value of g(2) is larger than the value of g(-4). The value of g(2) is smaller than the value of g(-4). The value of g(2) is the same as the value of g(-4). The values of g(2) and g(-4) cannot be compared.

OpenStudy (anonymous):

B the value of g(2) is smaller than the value of g(-4)

OpenStudy (mathmale):

Really, it's great that you're sharing your reason(s) for selecting choice B. Provided that your values for g(2) and g(-4) are correct, your overall answer is also correct. Nice work.

OpenStudy (anonymous):

can you help with3 more?